OpenStudy (erikaxx):

can someone help me with this one.... eliminate the arbitary constant of x^2 + y^2 - Ay =B the answer is xy`` - (y`)^3 - y` = 0

OpenStudy (erikaxx):

\[equation : x ^{2} + y ^{2} -Ay = B\]

OpenStudy (erikaxx):

\[final answer: xy ^{``} - (y ^{`})^{3} - y ^{`} = 0\]

OpenStudy (erikaxx):

@rvc

OpenStudy (erikaxx):

@SamsungFanBoy @zzr0ck3r @across

OpenStudy (erikaxx):

@baru

OpenStudy (baru):

can you differentiate the given equation?

OpenStudy (erikaxx):

\[x^2 + y^2 - Ay = B\] @baru

OpenStudy (baru):

yes, differentiate that with respect to x

OpenStudy (erikaxx):

how?

OpenStudy (baru):

\[\frac{d(x^2 + y^2 - Ay}{dx} = \frac{d (B)}{dx}\]

OpenStudy (baru):

\[2x + 2yy'-Ay' = 0\]

OpenStudy (erikaxx):

ah. thats what you mean by differentiate

OpenStudy (erikaxx):

okay got that part

OpenStudy (baru):

\[2x + 2yy'-Ay' = 0\] can you differentiate this again?

OpenStudy (erikaxx):

\[d(2x + 2yy`- Ay`)/dx = 0\]

OpenStudy (baru):

yes, what so what do you get after you carry out the differentiation

OpenStudy (erikaxx):

\[2 + y` - A = 0\]

OpenStudy (baru):

no, y is a function of x you have to differentiate with respect to x \[dy/dx=y'\\d(y')/dx=y''\]

OpenStudy (erikaxx):

ah, so 2y + 2yy`` - A yy`

OpenStudy (baru):

\[d(2x + 2yy`- Ay`)/dx = 0 \\=2 + 2yy'' +2y'y' - Ay''\]

OpenStudy (erikaxx):

why Ay` become 2y`y1?

OpenStudy (erikaxx):

no.. i mean why theres a 2y`y`?

OpenStudy (baru):

its called the 'product rule' of differentiation\[\frac{d(2yy')}{dx} = 2y \frac{d(y')}{dx} + y' \frac{d(2y)}{dx}\]

OpenStudy (erikaxx):

didnt know about that. whats the formula for that?

OpenStudy (baru):

if u and v are functions of x\[\frac{d(uv)}{dx}=u \frac{dv}{dx} + v \frac{du}{dx}\]

OpenStudy (erikaxx):

ah, so you use the power rule on 2yy`?

OpenStudy (baru):

i used the 'product rule'

OpenStudy (erikaxx):

yes the product rule i mean

OpenStudy (erikaxx):

so what next

OpenStudy (baru):

\[x^2 + y^2 - Ay = B\\ 2x + 2yy'-Ay' = 0\\ 2 + 2yy'' +2y'y' - Ay''=0\] now we need you use these three equations to some how get rid of A and B

OpenStudy (baru):

re-arrange the second equation \[2x + 2yy'-Ay' = 0\\2x + 2yy' = Ay'\\A=\frac{2x}{y'} + 2y\]

OpenStudy (baru):

substitute A in the third equation

OpenStudy (baru):

\[2 + 2yy'' +2y'y' - ( \frac{2x}{y'}+2y)y''=0\]

OpenStudy (erikaxx):

yes i got the same thing

OpenStudy (erikaxx):

so y` and the y in 2y will be cancel?

OpenStudy (baru):

which part are you talking about

OpenStudy (erikaxx):

the substituting the A

OpenStudy (baru):

i just divided both sides by y'

OpenStudy (erikaxx):

ah, okay got that one

OpenStudy (erikaxx):

so will going to subtract all the equations?

OpenStudy (baru):

no, we substituted A and thus eliminated it \[2 + 2yy'' +2y'y' - ( \frac{2x}{y'}+2y)y''=0\] all thats left is to simplify the above equation so that it matches the given answer

OpenStudy (erikaxx):

but the answer is \[xy` -(y`)^3 - y` = 0\]

OpenStudy (baru):

as i said, we need to simplify the above equation so just multiply both sides by \(\frac{y'}{2}\)

OpenStudy (baru):

do that and you should have your answer

OpenStudy (erikaxx):

all the side by 2 so.. \[yy` +y`y` (x/y` +y) y` = 0\]

OpenStudy (baru):

\[(2 + 2yy'' +2y'y' - ( \frac{2x}{y'}+2y)y'')) \times \frac{y'}{2}=0 \times \frac{y'}{2}\]

OpenStudy (erikaxx):

why u didnt divide the whole equation by 2?

OpenStudy (erikaxx):

so y` + y`` + y` - x y`

OpenStudy (baru):

i get \[y' + yy'y'' +(y')^3 -xy''-yy'y''=0\] simplify further \[y' +(y')^3 -xy''=0\] multiply both sides by -1 and rearrange \[xy'' - (y')^3 -y' =0\]

OpenStudy (erikaxx):

ah wait why you multipley all sides by y`/2?

OpenStudy (baru):

i did that to simplify the expression